Prove that the vector A = 3yz2 i + 4x322 j - 3x2y2 k is solenoidal


1
Expert's answer
2021-03-03T05:53:38-0500

A is solenoidal if "\nabla . A = 0"

"\nabla. A = \frac{\partial(3yz^2)}{\partial x} + \frac{\partial(4x^3z^2)}{\partial y} - \frac{\partial(3x^2y^2)}{\partial z} = 0+0-0 = 0"

as desired.